OpenStudy (anonymous):

How do I write the linear equation in slope - intercept form with slope=2/3 and y- intercept =-1

OpenStudy (neonumbrella5115):

Do you know what slope intercept form is?

OpenStudy (anonymous):

Yea

OpenStudy (neonumbrella5115):

Can you type it out? Because after that, we will plug the slope in for m and the y intercept for b.

OpenStudy (anonymous):

Slope= 2\3 yintercept= -1

OpenStudy (neonumbrella5115):

That's what each of those values equal but not the actual form of the equation we're looking form. Can you try again?

OpenStudy (anonymous):

Its slope- intercept Y= MX +B

OpenStudy (mathmale):

True. But... we were asking you for the EQUATION of a strai8ght line in slope-intercept form. Point slope form is y-y1 = m (x - x1). This is not what we wanted, but it does illustrate one form of the equation of a line. Slope-intercept form? You research this... type out the proper form Essential that you know this form of the equation of a straight line.

OpenStudy (neonumbrella5115):

So now that you know that, you can plug in the slope (2/3) for the m value. Then plug in the y intercept (-1) for the B.

OpenStudy (anonymous):

So it would be y= 2\3 -1

OpenStudy (neonumbrella5115):

Yes! Good job! That is the correct answer!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

So what about 2x+8y=16

OpenStudy (tkhunny):

Well, with an 'x' in there, anyway.

OpenStudy (tkhunny):

Jus solve it for y. Show your best algebra skills.

OpenStudy (neonumbrella5115):

@tkhunny Yes that too. So for solve for y, and then once you do that, the equation will end up in slope intercept form

OpenStudy (anonymous):

Would I subtract the 8y from 2y

OpenStudy (neonumbrella5115):

When? Can you post the steps you are going through?

OpenStudy (anonymous):

Yes 2x+8y=16

OpenStudy (anonymous):

That's all I got

OpenStudy (neonumbrella5115):

So first you would subtract 2x from both sides.

OpenStudy (anonymous):

2x-8x=16 -2x. -2x

OpenStudy (tkhunny):

Let's keep track of the "8y", shall we?

OpenStudy (mathmale):

2x-2x is zero. That leaves you with -8y on the left side. On the right side you'll have 16-2x. Write out this new form of the equation. Then, consider how you'd solve for y.